OpenStudy (anonymous):

solve. Square root of x + 9 = x - 3 Show work.

OpenStudy (jdoe0001):

\(\bf \sqrt{x + 9} = x - 3\implies (\sqrt{x + 9})^2=(x - 3)^2\implies x+9=(x - 3)^2 \\ \quad \\ x+9=(x - 3)(x - 3)\) expand and solve the quadratic

OpenStudy (anonymous):

I wish it was that easy but the only answer choices are in fraction form. @jdoe0001

OpenStudy (jdoe0001):

well..... I"d assume you'd end up using the quadratic formula, which I'd think you've covered by now then